# On the Convergence of Hermitian Dynamic Mode Decomposition

### v1

@@ 4 A general convergence result
Throughout this section, we consider an arbitrary self-adjoint operator $\mathcal{L}$ acting on a Hilbert space $\mathcal{H}$ with domain $\mathcal{D}(\mathcal{L})$. Let $\{\mathcal{P}_{n}:\mathcal{H}\rightarrow\mathcal{V}_{n}\}_{n\in\mathbb{N}}$ be a sequence of orthogonal projections onto a Hilbert space $\mathcal{V}_{n}\subset\mathcal{H}$, such that $\mathcal{P}_{n}^{*}\mathcal{P}_{n}$ converges strongly to the identity, meaning that
@@ 4.1 Bounded operators - the easy case
The above proof cannot be carried over to unbounded operators since the moments may not exist. Instead, we replace the polynomials appearing in the proof of Lemma 4.1 with rational functions. To do this we must look at the resolvent.
@@ 4.2 Strong convergence of the resolvent
Without such assumptions, spectral measures need not converge (Levitin, , Theorem 2.3). The following lemma shows that with this assumption, the resolvent of our projected operators converge strongly to the resolvent of $\mathcal{L}$. For a discussion of these kinds of results, see (kato2013perturbation, , Chapter VIII) and weidmann1997strong ; rellich1941storungstheorie .
@@ 4.3 General convergence theorem
The convergence of Hermitian DMD now follows almost immediately. Then,for a given observable $g\in L^{2}(\Omega,\omega)$, we define
@@
$\lim_{M\rightarrow\infty}\mathbf{g}_{N,M}=\mathbf{g}_{N},\quad\text{with}\quad\mathcal{P}_{N}g=\Psi\mathbf{g}_{N}.$
We emphasize that this is the standard way to compute expansion coefficients in the EDMD algorithm.
@@
were $\mu_{\mathbf{g}_{N,M}}^{(M)}$ is the spectral measure of $\Psi\mathbf{g}_{N,M}$ with respect to the Hermitian DMD matrix.
@@ 5 Numerical example
$i\frac{\partial\Psi}{\partial t}=\hat{H}(\Psi)=-\frac{1}{2}\Delta\Psi+V(x,y)\Psi,\quad(x,y)\in\mathbb{R}^{2},$
where $V(x,y)=(x^{2}+y^{2})/2$ is an external harmonic potential, $\Psi:\Omega\to\mathbb{C}$ is a normalized wave function, and $\hat{H}$ is the Hamiltonian operator of the system. Here, we select a computational domain $\Omega=(-5,5)^{2}$ to be large enough so that solutions vanish well before reaching the boundary. The snapshot data contain functions of the form $(\Psi,i\partial_{t}\Psi)$, which are related by the self-adjoint Hamiltonian operator $\hat{H}$. We use $N=400$ initial conditions consisting of Gaussian bumps (the dictionary) inside $\Omega$ as
$\Psi(x,y)=(1+i)e^{-3((x-x_{i})^{2}+(y-y_{j})^{2})},\quad 1\leq i,j\leq 20,$
where the Gaussian centers $(x_{i},y_{j})$ are uniformly distributed in the domain $[-4,4]^{2}$. We then evaluate the functions on the domain $\Omega$ at a uniform tensor-product grid of $M=300^{2}$ snapshot points, and discretize the integral in (2.2) using a trapezoidal rule. Finally, we compute the $N\times N$ Koopman matrix $\mathbf{K}$ using the Hermitian DMD algorithm described in Section 3, and perform an eigenvalue decomposition of $\mathbf{K}$ to obtain the first eigenstates and corresponding eigenvalues to the Schrödinger operator (see Fig. 2).
@@
Figure 2: The first $6$ energy eigenstates $|\psi_{j}|^{2}$ and eigenvalues of the Schrödinger operator discovered by Hermitian DMD.
The analytical expression for the eigenmodes of the Schrödinger equation are obtained using the standard ansatz $\Psi(x,y,t)=\psi(x,y)e^{-iEt}$, where $\phi$ is the eigenfunction and $E$ is the corresponding eigenvalue. Following a separation of variables, one finds that the eigenmodes are proportional to (charalampidis2018computing, , Eq. (8))
$\psi_{m,n}(x,y)=H_{m}(x)H_{n}(y)e^{-(x^{2}+y^{2})/2},\quad(x,y)\in\mathbb{R}^{2},\quad m,n\geq 0,$
@@
Here, $(\lambda_{j},\mathbf{v}_{j})$ are the eigenpairs computed by Hermitian DMD and $f=\Psi\mathbf{f}$. For comparison, for each analytic eigenvalue, we take the mean of the cluster of the eigenvalues $\lambda_{j}$ that approximate it and then sum the weights $c_{j}$. The results are shown in Figs. 4 and 1 and demonstrate the convergence in Theorem 4.6.
@@
Table 1: Convergence of the $c_{j}$ ($\lambda_{j}$) in the first five spikes of Fig. 4 for different values of $N$.

### v2

@@ 4 A general convergence result
Throughout this section, we consider an arbitrary self-adjoint operator $\mathcal{L}$ acting on a Hilbert space $\mathcal{H}$ with domain $\mathcal{D}(\mathcal{L})$. For example, the differential operator $-\mathrm{d}^{2}/\mathrm{d}x^{2}$ on $L^{2}(\mathbb{R})$ is self-adjoint with domain $\{g\in L^{2}(\mathbb{R}):g^{\prime\prime}\in L^{2}(\mathbb{R})\}$. This is an example of an unbounded operator (recall that an operator $A$ is bounded if $\sup_{x\in\mathcal{D}(A),\|x\|=1}\|Ax\|<\infty$). Moreover, the operators in this section need not be Koopman operators. Let $\{\mathcal{P}_{n}:\mathcal{H}\rightarrow\mathcal{V}_{n}\}_{n\in\mathbb{N}}$ be a sequence of orthogonal projections onto a Hilbert space $\mathcal{V}_{n}\subset\mathcal{H}$, such that $\mathcal{P}_{n}^{*}\mathcal{P}_{n}$ converges strongly to the identity, meaning that
@@ 4.1 Bounded operators - the easy case
The above proof cannot be carried over to unbounded operators since the moments may not exist. Instead, we replace the polynomials appearing in the proof of Lemma 4.1 with rational functions. To do this, we must look at the resolvent.
@@ 4.2 Strong convergence of the resolvent
Without such assumptions, spectral measures need not converge [48, Theorem 2.3]. The following lemma shows that with this assumption, the resolvents of our projected operators converge strongly to the resolvent of $\mathcal{L}$. For a discussion of these kinds of results, see [49, Chapter VIII] and [50, 51].
@@ 4.3 General convergence theorem
The convergence of Hermitian DMD now follows almost immediately. For a given observable $g\in L^{2}(\Omega,\omega)$, we define
@@
$\lim_{M\rightarrow\infty}\mathbf{g}_{N,M}=\mathbf{g}_{N},\quad\text{with}\quad\mathcal{P}_{N}g=\mathbf{\Psi}\mathbf{g}_{N},$
where $\mathbf{\Psi}$ is the dictionary feature map in Eq. 2.2. We emphasize that this is the standard way to compute expansion coefficients in the EDMD algorithm.
@@
were $\mu_{\mathbf{g}_{N,M}}^{(M)}$ is the spectral measure of $\mathbf{\Psi}\mathbf{g}_{N,M}$ with respect to the Hermitian DMD matrix.
@@ 5 Numerical example
$i\frac{\partial u}{\partial t}=\hat{H}(u)=-\frac{1}{2}\Delta u+V(x,y)u,\quad(x,y)\in\mathbb{R}^{2},$
where $V(x,y)=(x^{2}+y^{2})/2$ is an external harmonic potential, $u:\Omega\to\mathbb{C}$ is a normalized wave function, and $\hat{H}$ is the Hamiltonian operator of the system. Here, we select a computational domain $\Omega=(-5,5)^{2}$ to be large enough so that solutions vanish well before reaching the boundary. The snapshot data contain functions of the form $(u,i\partial_{t}u)$, which are related by the self-adjoint Hamiltonian operator $\hat{H}$. We use $N=400$ initial conditions consisting of Gaussian bumps (the dictionary) inside $\Omega$ as
$u(x,y)=(1+i)e^{-3((x-x_{i})^{2}+(y-y_{j})^{2})},\quad 1\leq i,j\leq 20,$
where the Gaussian centers $(x_{i},y_{j})$ are uniformly distributed in the domain $[-4,4]^{2}$. [^4] For suitable Gaussian bump functions, one can easily show that Assumption 4.3 is satisfied for our Schrödinger operator. We then evaluate the functions on the domain $\Omega$ at a uniform tensor-product grid of $M=300^{2}$ snapshot points, and discretize the integral in (2.3) using a trapezoidal rule. Finally, we compute the $N\times N$ Koopman matrix $\mathbf{K}$ using the Hermitian DMD algorithm described in Section 3, and perform an eigenvalue decomposition of $\mathbf{K}$ to obtain the first eigenstates and corresponding eigenvalues to the Schrödinger operator (see Fig. 2).
@@
Figure 2: The first $6$ energy eigenstates and eigenvalues of the Schrödinger operator discovered by Hermitian DMD.
The analytical expression for the eigenmodes of the Schrödinger equation are obtained using the standard ansatz $u(x,y,t)=\phi(x,y)e^{-iEt}$, where $\phi$ is the eigenfunction and $E$ is the corresponding eigenvalue. Following a separation of variables, one finds that the eigenmodes are proportional to [52, Eq. (8)]
$\phi_{m,n}(x,y)=H_{m}(x)H_{n}(y)e^{-(x^{2}+y^{2})/2},\quad(x,y)\in\mathbb{R}^{2},\quad m,n\geq 0,$
@@
Here, $(\lambda_{j},\mathbf{v}_{j})$ are the eigenpairs computed by Hermitian DMD and $f=\mathbf{\Psi}\mathbf{f}$. For comparison, for each analytic eigenvalue, we take the mean of the cluster of the eigenvalues $\lambda_{j}$ that approximate it and then sum the weights $c_{j}$. The results are shown in Figs. 4 and 1 and demonstrate the convergence in Theorem 4.6.
Table 1: Convergence of $c_{j}$ ($\lambda_{j}$) in the first five spikes of Fig. 4 for different values of $N$.
@@
Data availability
The code to reproduce the numerical experiments is publicly available on GitHub at https://github.com/NBoulle/ConvergenceHermitianDMD.
Acknowledgments
NB was supported by the SciAI Center funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729.